Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 57     Entry time: 20 Mar 2024, 23:03
Author: Giuseppe Mercurio 
Group: SCS 
Subject: Shift summary 

Late shift summary

  • coarse timing
  • installed laser shutter on the laser breadboard and I0 diode
  • exchanged camera of microscope to improve resolution
  • knife edge scans of FEL after BOZ
  • realized that knife edge scans of OL cannot be done with present setup because OL cannot reach the diode in DIAG
  • Nahid prepared FeByP 1 liter, 30 mM
  • preparation of liquid jet for the night

problems:

*BOZ online analysis not yet working (in progress)
*OL cannot reach diode in CHEM_DIAG
*one lasermet shutter in ILH is not operational, we use the one meant for the XOX branch
*new ADC karabo device stops frequently, so we went back to the normal FADC2 device to be able to do knife edge scans. The scenes are linked to the main CHEM scene.

ELOG V3.1.4-7c3fd00